Rebekah's recipe for 4 1/2 dozen cookies calls for 6 cups of sugar. How many cups of sugar are needed to make 8 dozen cookies?
Please help
Thanks

Answer by checkley77(12844)

54/6=96/x
54x=96*6
54x=576
x=576/54
x=10.6667 cups of sugar are required.